### v1

# arXiv:2603.03220v1[cond-mat.mtrl-sci] 03 Mar 2026

Abstract
A novel implementation of the linear response time-dependent density functional theory addressing spin excitations in non-collinear magnets based on the Korringa-Kohn-Rostoker Green’s function method is presented. Following the exposition of the formalism based on the adiabatic local spin density approximation to the exchange-correlation kernel generalized to the non-collinear case, the computational scheme is discussed in detail. The formation of the Goldstone modes in non-collinear susceptibility calculations is elaborated on formally and from the numerical convergence point of view. The scheme is deployed to study the dispersion, Landau damping, and spatial shapes of magnons for the representative members of the kagome non-collinear antiferromagnets.
1 Introduction
The ab initio linear response time-dependent density functional theory (LRTDDFT) of spin excitations has become an indispensable tool in the computational design of novel functional materials [Gross1985, Savrasov1998, Staunton1999, Buczek2009, Durhuus2023, Binci2025]. It offers a comprehensive description not only of spin-waves (magnons), which can be also studied by means of an empirical mapping onto the celebrated Heisenberg model [LeBlanc2021], but also of Stoner excitations and of the associated Landau damping, inherent to the itinerant magnetism [Moriya1985]. No doubt the dire need of the moment is the extension of the LRTDDFT to non-collinear (NC) magnets, i.e., spin-polarized solids where the direction of magnetization is allowed to vary as a function of position [Sandratskii1998]. In the spintronic and magnonic context, an enormous practical potential has been ascertained to the members of this material family [1, Rimmler2024]. Moreover, they often accompany other exotic or strongly correlated phases of matter, e.g., the superconductivity [Bruening2025]. However, the true attractive force of these magnets stems from their beautiful symmetries [Sandratskii1991] rooted deeply in the modern field theories [Azaria1993, Pradenas2024, Pradenas2025]
A pivotal role in the development of magnonic devices is played by the damping of the spin-waves [Chumak2015, Joshi2016, Duine2018]. Except for the Landau mechanism, spin-waves can also scatter considerably on crystal imperfections [Buczek2016]. Remarkably, NC magnets cannot be fully grasped within the linear spin-wave theory and their magnons undergo intrinsic damping due to two-boson interactions even at absolute zero [Chernyshev2009, Zhitomirsky2013, Park2020] contrary to the case of collinear ferromagnets [Paischer2024].
Prompted to close this important methodological gap, in this paper, we present the first computer implementation of the LRTDDFT for NC magnets and deploy it to study collective spin excitations in representative members of the NC kagome antiferromagnets (KAF) [Park2018, Park2020].
The paper is organized as follows. Sections 2 and 3 expose the underlying formalism and our implementation based on the Korringa-Kohn-Rostoker Green’s function method. In section 4, we discuss the dispersion, Landau damping, and spatial shapes of magnons for the representative members of the NC KAF.
2 Formalism
2.1 Generalized susceptibility
The following exposition pertains to a general interacting electron system with a possibly NC ground state. Unless stated otherwise, Rydberg atomic units are used throughout, with $\hbar=1$, Bohr radius $a_{0}=1$, and Rydberg energy $E_{\mathrm{R}}=1$ which in turn implies the numerical values of the electron charge and mass to be respectively $e=\sqrt{2}$ and $m_{e}=\frac{1}{2}$.
Here, the description of properties of spin and charge excitations is facilitated by the evaluation of the retarded density-density response function [Fetter1971]
$\chi^{ij}{\left({\mathbf{x}},{\mathbf{x}}^{\prime},t-t^{\prime}\right)}$ $=-{{\mathfrak{i}}}{\theta}(t-t^{\prime}){\left\langle{{\left[\hat{\sigma}^{i}{\left({\mathbf{x}}t\right)},\hat{\sigma}^{j}{\left({\mathbf{x}}^{\prime}t^{\prime}\right)}\right]}}\right\rangle}.$ (1)
In the linear regime, the quantity yields the charge or magnetization density response $\delta n^{i}{\left({\mathbf{x}},t\right)}$ emerging due to the dynamical perturbing fields $\delta V^{j}{\left({\mathbf{x}}^{\prime},t^{\prime}\right)}$ (magnetic or scalar) coupling to these densities. $\hat{\sigma}^{i}{\left({\mathbf{x}}t\right)}$ are charge ($i=0$) and magnetization density operators ($i=x,y,z$), ${{\left[A,B\right]}}\equiv AB-BA$ is the commutator, and the $^{\left\langle\hat{o}\right\rangle}$ is the expectation value of the operator $\hat{o}$ for the system without the perturbation. Throughout all the paper, $\mathbf{x}$} refers to the global position vector, whereas $\mathbf{r}$} is the position within the primitive unit cell. It is assumed that the potentials of the unperturbed system bear no time dependence in the Schrödinger picture. Note that the response function can be used to study any electron density excitations including magnons, excitons, and plasmons.
The systems studied in this work are assumed to feature translational invariance which makes the response functions diagonal in the Bloch basis
$\chi{\left({\mathbf{r}},{\mathbf{r}}^{\prime},{\mathbf{q}}\right)}=\sum_{{\mathbf{R}}\in\mathcal{L}}{\mathfrak{e}^{-{{\mathfrak{i}}}{\mathbf{q}}\cdot{\mathbf{R}}}}\chi{\left({\mathbf{r}}+{\mathbf{R}},{\mathbf{r}}^{\prime}\right)}.$ (2)
Here, $\mathbf{r}$} and $\mathbf{r}$$^{\prime}$ belong to the same primitive cell of the crystal and $\mathcal{L}$ denotes the set of lattice vectors. (In the equation above other arguments are suppressed for simplicity.)
In the frequency domain, the anti-Hermitian part of the susceptibility (which we refer to as a loss matrix)
$L^{ij}{\left({\mathbf{r}},{\mathbf{r}}^{\prime},{\mathbf{q}},\omega\right)}\equiv\frac{1}{2{{\mathfrak{i}}}}{\left(\chi^{ij}{\left({\mathbf{r}},{\mathbf{r}}^{\prime},{\mathbf{q}},\omega\right)}-\chi^{ji}{\left({\mathbf{r}}^{\prime},{\mathbf{r}},{\mathbf{q}},\omega\right)}^{\ast}\right)}.$ (3)
allows to analyse the spectra and the spatial forms of the excited density modes. For a specific energy $\omega$, the eigenvectors $\xi_{\lambda}^{i}{\left({\mathbf{r}},{\mathbf{q}},\omega\right)}$ of $L$ represent the shapes of the natural modes of charge and magnetization density fluctuations at this frequency. The magnitude of the associated eigenvalues, $a_{\lambda}{\left({\mathbf{q}},\omega\right)}$, gives the density of states of these modes, up to a factor $1/\pi$. In other words, large $a_{\lambda}$ signifies the presence of collective density modes of the system. These modes are real physical objects of the following form
${\mathfrak{Re}{\left[\xi_{\lambda}^{i}{\left({\mathbf{r}},{\mathbf{q}},\omega\right)}{\mathfrak{e}^{{{\mathfrak{i}}}{\left({\mathbf{q}}\cdot{\mathbf{R}}-\omega t-\phi\right)}}}\right]}},$ (4)
where $\phi$ is an overall phase of the mode. Following the fluctuation dissipation theorem, which we briefly outline in Appendix A, the loss matrix has several further useful physical interpretations. On one hand, $\omega a_{\lambda}{\left({\mathbf{q}},\omega\right)}/2$ yields the power absorbed by the system from the driving external fields of the shape $\xi_{\lambda}^{i}{\left({\mathbf{r}},{\mathbf{q}},\omega\right)}$ coupling to the densities. On the other, it can be used to compute the expectation values of the time correlations ${\mathfrak{C}}^{ij}{\left({\mathbf{r}},{\mathbf{r}}^{\prime},{\mathbf{q}},\omega\right)}$ between density operators in the unperturbed state, i.e., among others, the spontaneous fluctuations of the system density.
Furthermore, the magnetization density loss matrix evaluated in the momentum space is proportional to the cross-section of the inelastic neutron scattering off spin-density excitation [VanHove1954]. Last but not least, the quantity is related to the (spin-polarized) electron energy loss spectroscopy ((SP)EELS) signal [Kirschner1985, Vollmer2003, Costa2006, Santos2018]. In the latter case, however, contrary to the neutron scattering, the loss energy dependence of the cross-section involves intricate details of the Coulomb interaction between the electrons of the beam and the probe [Vignale1985, Plihal1998, Hong2000a, Paischer2024b].
2.2 Linear response time dependent density function theory formulation
The evaluation of the dynamic susceptibilities of interacting electron gas is as involved as the solution of the many-body problem itself [Abrikosov1975] and one must resort to approximations. A successful ab initio scheme can be derived within the LRTDDFT [Gross1985, Qian2002] and it often involves two steps. First, the so called Kohn-Sham susceptibility, $\chi_{\mathrm{KS}}$ is constructed describing the response of the formally non-interacting Kohn-Sham (KS) system to an external perturbation. Second, the self-consistent dynamical change in the exchange-correlation potential arising due to the densities induced by the external field and this change is incorporated upon solving so called susceptibility Dyson equation. There are different complementary approaches to implement this scheme and some do not explicitly break it into these two distinct steps [Savrasov1998, Binci2025]. The magnetic response function can be also found within the many-body perturbation scheme [Aryasetiawan1999, Karlsson2000, Mueller2016, Nabok2021] and using empirical Hubbard-type Hamiltonians [Cooke1980, Gumbs1980, Tang1998, Muniz2002, Bonetti2022].
Here, we discuss the form of the formalism underlying our implementation. The KS system is described by a corresponding formally single-electron Green’s function (GF)
$G_{\alpha\beta}{\left({\mathbf{x}},{\mathbf{x}}^{\prime},z\right)}=\sum_{j}\frac{\phi_{j}{\left({\mathbf{x}}\alpha\right)}\phi_{j}{\left({\mathbf{x}}^{\prime}\beta\right)}^{\ast}}{z-{\epsilon}_{j}},\,z\in{\mathbb{C}},$ (5)
where $\phi_{j}{\left({\mathbf{x}}\alpha\right)}$’s and ${\epsilon}_{j}$’s stand for KS eigenfunctions and eigenenergies, respectively, and $\alpha,\beta,\ldots$ denote spin degrees of freedom (up or down along a chosen quantization axis). The charge and magnetization densities, necessary for the Kohn-Sham potential’s construction, can be found by complex integration of $G$.
In the Korringa-Kohn-Rostoker method [Mankovsky2011, Simon2019], the Green’s function is not constructed from this Källén-Lehmann spectral representation but found directly upon solving the Dyson equation, written formally as
$G{\left(z\right)}=G_{0}{\left(z\right)}+G_{0}{\left(z\right)}V{\left(z\right)}G{\left(z\right)}.$ (6)
(For simplicity the arguments of the Green’s functions as well as corresponding sums and integrals are suppressed.) The ground state potential $V$ includes the static external fields as well as the ground state Hartree and exchange-correlation contributions. In this case it does not feature the dependence on the complex energy $z$ and is local. This is the form used to describe the materials considered in this paper.
Additionally, in many cases, it is necessary to include terms originating from the spin-orbit coupling (SOC) [Paischer2024]. Furthermore, the frequency dependence arises when typical electronic self-energy terms are incorporated, as it is the case when the disorder is described in the coherent potential approximation (CPA) [Buczek2016, Paischer2021, Paischer2021a] or many-body effects like magnon-electron interaction are included [Paischer2023, Usachov2024, Paischer2024a]. In the latter case the self-energy $V$ can additionally become non-local.
With the KS GF at hand, one can apply the Cauchy theorem in order to obtain the retarded KS susceptibility analytically continued to a line in a finite distance $\gamma$ from the real axis (the so called nearly real axis) [Abrikosov1975, Schmalian1996] ($\gamma>\gamma^{\prime}={0^{+}}$):
$\chi^{ij}_{\mathrm{KS}}$ ${\left({\mathbf{x}},{\mathbf{x}}^{\prime},\omega+{{\mathfrak{i}}}\gamma\right)}=-\frac{1}{2\pi{{\mathfrak{i}}}}\int_{-\infty}^{\infty}d{\epsilon}\times{}$
$($ ${f_{T}}{\left({\epsilon}+{{\mathfrak{i}}}\gamma^{\prime}\right)}S_{ij}{\left({\mathbf{x}},{\mathbf{x}}^{\prime},{\epsilon}+\omega+{{\mathfrak{i}}}\gamma+{{\mathfrak{i}}}\gamma^{\prime},{\epsilon}+{{\mathfrak{i}}}\gamma^{\prime}\right)}{}$
- ${f_{T}}{\left({\epsilon}-{{\mathfrak{i}}}\gamma^{\prime}\right)}S_{ij}{\left({\mathbf{x}},{\mathbf{x}}^{\prime},{\epsilon}+\omega+{{\mathfrak{i}}}\gamma-{{\mathfrak{i}}}\gamma^{\prime},{\epsilon}-{{\mathfrak{i}}}\gamma^{\prime}\right)}{}$
$+$ ${f_{T}}{\left({\epsilon}+{{\mathfrak{i}}}\gamma^{\prime}\right)}S_{ij}{\left({\mathbf{x}},{\mathbf{x}}^{\prime},{\epsilon}+{{\mathfrak{i}}}\gamma^{\prime},{\epsilon}-\omega-{{\mathfrak{i}}}\gamma+{{\mathfrak{i}}}\gamma^{\prime}\right)}{}$
- ${f_{T}}{\left({\epsilon}-{{\mathfrak{i}}}\gamma^{\prime}\right)}S_{ij}{\left({\mathbf{x}},{\mathbf{x}}^{\prime},{\epsilon}-{{\mathfrak{i}}}\gamma^{\prime},{\epsilon}-\omega-{{\mathfrak{i}}}\gamma-{{\mathfrak{i}}}\gamma^{\prime}\right)})$ (7)
where the $S$ is obtained as a trace of the product of two GFs and Pauli matrices $\sigma^{i}$.
$S_{ij}{\left({\mathbf{x}},{\mathbf{x}}^{\prime},z_{1},z_{2}\right)}\equiv\sigma^{i}_{\alpha\beta}G_{\beta\gamma}{\left({\mathbf{x}},{\mathbf{x}}^{\prime},z_{1}\right)}\sigma^{j}_{\gamma\delta}G_{\delta\alpha}{\left({\mathbf{x}}^{\prime},{\mathbf{x}},z_{2}\right)}.$ (8)
${f_{T}}({\epsilon}_{j})$ stands for the Fermi-Dirac distribution function. A small finite temperature is introduced to smear slightly the Fermi level allowing for better convergence of the Brillouin zone integrals. In this work, this temperature is negligible on the electronic energy scales and the electron gas is regarded to be described at absolute zero. $0^{+}$} denotes the small positive energy limit. Einstein summation convention is assumed for the spin indices and there is no difference raised and lowered density index $i$. The advantage of this approach is the possibility of deforming the complex plane integration contours such that the evaluation of $S$ takes place far from the singularities of the KS Green’s function allowing for its more rapid numerical convergence [Buczek2011a]. Furthermore, the evaluation of the susceptibility on the nearly real axis yields more stable analytic continuation to the real frequencies than the evaluation at Matsubara poles [Staunton1999, Staunton2000, Thakor2003].
The susceptibility Dyson equation allows to find the true interacting (enhanced) susceptibility of the many-electron system:
$\chi^{ij}$ ${\left({\mathbf{x}},{\mathbf{x}}^{\prime},\omega\right)}=\chi^{ij}_{\mathrm{KS}}{\left({\mathbf{x}},{\mathbf{x}}^{\prime},\omega\right)}+$ (9)
$\sum_{k,l=0}^{3}\iint d{\mathbf{x}}_{1}d{\mathbf{x}}_{2}\chi^{ik}_{\mathrm{KS}}{\left({\mathbf{x}},{\mathbf{x}}_{1},\omega\right)}{\left(\frac{2\delta_{k0}\delta_{l0}}{{\left|{\mathbf{x}}_{1}-{\mathbf{x}}_{2}\right|}}+K_{\mathrm{xc}}^{kl}{\left({\mathbf{x}}_{1},{\mathbf{x}}_{2},\omega\right)}\right)}\chi^{lj}{\left({\mathbf{x}}_{2},{\mathbf{x}}^{\prime},\omega\right)}.$
The first term in the bracket above denotes the Hartree (Coulomb) interaction ($v_{\textrm{C}}{\left({\mathbf{x}}\right)}\equiv 2/x$) while $K_{\mathrm{xc}}$ stands for the exchange-correlation kernel, defined as a functional derivative of the exchange-correlation potential evaluated at the ground state values of electronic and magnetic densities
$K_{\mathrm{xc}}^{ij}{\left[{\left\langle\hat{{\boldsymbol{\sigma}}}{\left({\mathbf{x}}\right)}\right\rangle}\right]}{\left({\mathbf{x}},{\mathbf{x}}^{\prime},t-t^{\prime}\right)}\equiv\frac{\delta v_{\mathrm{xc}}^{i}{\left({\mathbf{x}},t\right)}}{\delta n^{j}{\left({\mathbf{x}}^{\prime}t^{\prime}\right)}}.$ (10)
2.3 Exchange-correlation kernel
The exact form of the $K_{\mathrm{xc}}$ is, in general, unknown, especially for NC systems. There is a steady progress in constructing NC spin density functionals [Sharma2007, Kurth2009, Eich2010, Eich2013, Eich2013a, Pittalis2017, Eich2018], but their their inclusion in practical LRTDDFT schemes is still in progress. In this work, we adopt the adiabatic local spin density approximation (ALSDA) yielding $K_{\mathrm{xc}}$ in the following form
$K_{\mathrm{xc}}^{ij}{\left[{\left\langle\hat{{\boldsymbol{\sigma}}}{\left({\mathbf{x}}\right)}\right\rangle}\right]}{\left({\mathbf{x}},{\mathbf{x}}^{\prime},t-t^{\prime}\right)}\approx\frac{\delta v_{\textrm{LSDA}}^{i}{\left[{\left\langle\hat{{\boldsymbol{\sigma}}}{\left({\mathbf{x}}\right)}\right\rangle},{\mathbf{x}}\right]}}{\delta n^{j}{\left({\mathbf{x}}\right)}}\delta{\left({\mathbf{x}}-{\mathbf{x}}^{\prime}\right)}\delta{\left(t-t^{\prime}\right)}.$ (11)
For NC systems, this kernel is most easily written in a local coordinate system with its $z$-axis oriented along the magnetization direction [vonBarth1972, Katsnelson2004]. Its space-diagonal part reads in this case
$K_{\mathrm{xc}}^{\textrm{loc}}=\begin{pmatrix}{\frac{\partial V_{\mathrm{xc}}}{\partial m}}&0&0&{\frac{\partial V_{\mathrm{xc}}}{\partial m}}\\ 0&K^{\perp}&0&0\\ 0&0&K^{\perp}&0\\ -{\mu_{\textrm{B}}}{\frac{\partial B_{\mathrm{xc}}}{\partial n}}&0&0&-{\mu_{\textrm{B}}}{\frac{\partial B_{\mathrm{xc}}}{\partial m}}\\ \end{pmatrix}$ (12)
The matrix above is indexed with $i,j=0,x,y,z$, the spatial argument $\mathbf{x}$} of the kernel has been omitted for brevity in $K_{\mathrm{xc}}^{\textrm{loc}}{\left({\mathbf{x}}\right)}$ and in the partial derivatives on the right hand side. The derivatives are taken at the ground state values of the densities. $V_{\mathrm{xc}}{\left({\mathbf{x}}\right)}$ and $B_{\mathrm{xc}}{\left({\mathbf{x}}\right)}$ are the local magnitudes of the exchange-correlation scalar potential and magnetic field. We note the symmetry
${\frac{\partial V_{\mathrm{xc}}}{\partial m}}=-{\mu_{\textrm{B}}}{\frac{\partial B_{\mathrm{xc}}}{\partial n}}$ (13)
and the fact that
$K^{\perp}{\left({\mathbf{x}}\right)}$ $=-{\mu_{\textrm{B}}}\frac{B^{\mathrm{GS}}_{\gamma}{\left({\mathbf{x}}\right)}}{m^{\mathrm{GS}}_{\gamma}{\left({\mathbf{x}}\right)}}$ (14)
$K^{\perp}$ is found simply as the ratio of the magnitudes of KS magnetic field and magnetization density. We recall that these two latter vectors are parallel in the LSDA. Furthermore, it is immediately evident that $K^{\perp}$ is isotropic in the subspace of the transverse magnetization direction. This form has a clear physical interpretation. Once a infinitesimal magnetization is induced in the direction perpendicular to ${\mathbf{m}}^{\mathrm{GS}}{\left({\mathbf{x}}\right)}$ the local ${\mathbf{B}}_{\mathrm{xc}}{\left({\mathbf{x}}\right)}$ rotates locally without changing its magnitude.
$K_{\mathrm{xc}}^{\textrm{loc}}$ is brought into the global coordinate system using a rotation matrix
$K_{\mathrm{xc}}=\mathcal{T}K_{\mathrm{xc}}^{\textrm{loc}}\mathcal{T}^{\dagger}$ (15)
given by
$\mathcal{T}=\begin{pmatrix}1&0&0&0\\ 0&\cos\vartheta\cos\varphi&-\sin\varphi&\sin\vartheta\cos\varphi\\ 0&\cos\vartheta\sin\varphi&\cos\varphi&\sin\vartheta\sin\varphi\\ 0&-\sin\vartheta&0&\cos\vartheta\end{pmatrix}$ (16)
with $\vartheta({\mathbf{x}})$ and $\varphi({\mathbf{x}})$ being the polar and azimuthal angles describing the local direction of the spin magnetization at point $\mathbf{x}$}. For simplicity, this spatial argument is omitted in the matrix above.
2.4 The emergence of the Goldstone modes in non-collinear magnets
The appearance of Nambu-Goldstone bosons (NGBs) modes due to the spontaneous breaking of continuous symmetries and the resulting presence of the degenerate manifold of ground-states is a well studied phenomenon [Nambu1960, Goldstone1961, Goldstone1962]. Magnetic orderings are its manifestations, too. In the case of collinear magnets, the original symmetry of the Hamiltonian, $\mathrm{SO}{\left(3\right)}$, spontaneously breaks down to $\mathrm{SO}{\left(2\right)}$ [Anderson1990, Peierls1991, Tscheuschner1992]. The broken symmetry generators are spin angular momentum components, $S_{i}$, the magnetic order parameter transform as a vector, and the manifold of vacua is the two-dimensional hypersphere.
However, the question of the actual number of NGBs and their dispersion in systems without Lorentz invariance featuring vacuum expectation values of field-theoretical charges is a more delicate one [Brauner2007, Watanabe2012]. Interestingly, as discussed by Nambu [Nambu2004], even the familiar cases of ferro- and antiferromagnetic orderings yield here two remarkably different pictures. In both cases there are two continuous broken symmetry generators, say $S_{x}$ and $S_{y}$, but in ferromagnets they become canonically conjugated yielding only one NGB mode with quadratic dispersion relation. On the contrary, in antiferromagnets, the two corresponding NGBs remain independent and both feature linear dispersion.
The spontaneous symmetry breaking in the case of a triangular antiferromagnet (TAF) is even more complex. The magnetic order parameter transforms as a three-dimensional body. As carefully discussed by Pradenas et al. [Pradenas2025], the full symmetry of the TAF Hamiltonian, $\mathrm{SO}{\left(3\right)}_{L}\times\mathrm{SO}{\left(2\right)}_{R,3}$, closely related to the one of a axially-symmetric rigid quantum top [Azaria1993], is spontaneously reduced to the residual symmetry $\mathrm{SO}{\left(2\right)}_{V,3}$. Three NGB branches with linear dispersion emerge [Chernyshev2009]. The residual symmetry leads to the conservation of the isospin and degeneracy of two spin-wave velocities. It is further broken down when long-range interactions between spins and the inter-plane magnetic coupling are taken into account, lifting the degeneracy of the two magnon branches [LeBlanc2014, LeBlanc2021].
While the emergence of the three Goldstone bosons in NC triangular ferromagnets can be be deduced from the general symmetry arguments outlined above, it is instructive to inquire how it manifests itself in the LRTDDFT calculations. Here, the correct long-wavelength behaviour, i.e., the presence of a mode or modes with vanishing energy for $q=0$, can be seen as a sum rule arising from the interplay between the exchange correlation kernel $K_{\mathrm{xc}}$ and the static uniform magnetic susceptibility [Lounis2010, Lounis2011, Buczek2011a, Mueller2016]. We name this interplay Goldstone sum rule and it can be stated as follows. The “Dyson denominator” of Eq. (2.2) for $q=0$ and $\omega=0$ (Coulomb interaction, $v_{\textrm{C}}$ plays no role in the formation of the magnetic NGB with anisotropy neglected)
${\sf I}-\chi_{\mathrm{KS}}K_{\mathrm{xc}}$ (17)
must feature zero eigenvalues. Their multiplicity, $N_{r}$, is equal to the number of linearly independent rigid rotations of the magnetic order parameter which, with spin-orbit coupling neglected, cost no energy. We refer to them as zero-energy magnetic deformations. They are the eigenvectors of the Dyson denominator with eigenvalue 0. In a NC magnets $N_{r}=3$ while $N_{r}=2$ holds for the collinear case, i.e., it is directly given by the number of broken symmetry generators. However, in the light of the preceding discussion, we note that $N_{r}$ is not necessarily equal to the number of Goldstone modes of a system.
Let us turn now to our specific case of the ALSDA. Apart from having no frequency dependence, this kernel, cf. Sec. 2.3, is $q$-independent and finite at $q=0$. The latter property tends to be a serious drawback in the charge-charge response channel, as it precludes the correct TDDFT description of excitonic effects in semiconductors [Onida2002, Turkowski2009]. On the other hand, it is the correct behaviour in the long wave-length limit for magnetization density channels, yielding correctly the magnon Goldstone mode which would not arise otherwise.
Generally speaking, there are two reasons why the Goldstone sum rule is not fulfilled exactly in numerical calculations.
The first is of systematic character and arises often in the many-body perturbation methods [Mueller2016, Nabok2021] where the ground state found self-consistently using an approach (typically LSDA to the ground state functional) different from the one for the kernel (corresponding to the screened Coulomb interaction $W$). Once $W$ is used to self-consistently find the magnetic ground state the deficiency will disappear. Let us briefly remark that the currently available many-body perturbation schemes for computing magnetic excitations are practically equivalent to the LRTDDFT in the ALSDA. They take advantage of the LSDA for the electronic ground state, local and static approximation to the screened Coulomb interaction $W$, as well as the choice of Feynman diagrams yielding the random phase approximation-like equation for the dynamic susceptibility resembling formally the equation (2.2) of the LRTDDFT.
The second reason are numerical inaccuracies arising during the computations of $\chi$. They can be broadly said to have two origins. The first includes the convergence of $\chi_{\mathrm{KS}}$ calculations, typically with respect to the Brillouin zone and energy integration as well as number of KS states used. Note that the latter aspect is of minor importance for the KKR-based implementations, contrary to the methods using direct Källén-Lehmann spectral representation for $\chi_{\mathrm{KS}}$ evaluation. The second origin of numerical errors is the finite or imperfect basis for the spatial representation of $\chi_{\mathrm{KS}}$ and $\chi$. In any case, accurate fulfilment of the Goldstone sum rule is crucial to the faithful description of magnon in the long wave-length regime and we look in detail now how it can be achieved in the case of the NC ground state and ALSDA.
Let us assume that $m^{\mathrm{GS}}_{\alpha a}$ and $B^{\mathrm{GS}}_{\alpha a}$ are ground state Kohn-Sham magnetic field and the magnetization density, respectively. Here, $\alpha$ encodes the vector direction and $a$ the spatial dependence of the quantities, i.e., the position, $a\equiv{\left({\mathbf{x}}\right)}$. The zero-energy magnetic deformation of the ground state can be written as
$m^{\beta}_{\alpha a}=\phi\sum_{\gamma}\epsilon_{\alpha\beta\gamma}m^{\mathrm{GS}}_{\gamma a}$ (18)
where $\epsilon_{\alpha\beta\gamma}\equiv{\left(\beta-\alpha\right)}{\left(\gamma-\alpha\right)}{\left(\gamma-\beta\right)}/2$ is the Levi-Civita symbol and $\phi$ an infinitesimal rotation angle of the ground state magnetization around the direction $\beta$. ($\alpha,\beta,\gamma\in\Set{1,2,3}\equiv\Set{x,y,z}$). Note that a linear combination of vectors $m^{\beta}_{\alpha a}$ corresponding to different $\beta$’s is a zero-energy magnetic deformation corresponding to an infinitesimal rotation around a specific axis as well. $m^{\beta}_{\alpha a}$ can be named “transverse” magnetic deformation because its direction (in the magnetization space) is orthogonal to the one of $m^{\mathrm{GS}}_{\alpha a}$. However, in the NC magnets, it does not lie in a single plane as in the collinear case.
When the deforming transverse magnetic field
$B^{\beta}_{\alpha a}=\phi\sum_{\gamma}\epsilon_{\alpha\beta\gamma}B^{\mathrm{GS}}_{\gamma a}$ (19)
is applied, the effective field of the Kohn-Sham system will rotate around $\beta$ by the angle $\phi$, too. When the Kohn-Sham equations are solved with this new field the resulting change of magnetization will become $m^{\beta}_{\alpha a}$ such that
$m^{\beta}_{\alpha a}=-{\mu_{\textrm{B}}}\sum_{\gamma b}\chi^{\mathrm{KS}}_{\alpha a\gamma b}{\left({\mathbf{q}}={\mathbf{0}},\omega=0\right)}B^{\beta}_{\gamma b}.$ (20)
$\chi^{\mathrm{KS}}_{\alpha a\gamma b}{\left({\mathbf{q}}={\mathbf{0}},\omega=0\right)}$ is a Hermitian matrix and thus features a complete set of eigenvectors. While the deformations $m^{\alpha}$ and $B^{\alpha}$ are clearly not eigenvectors of $\chi^{\mathrm{KS}}$, they can be decomposed in the eigenbasis of $\chi^{\mathrm{KS}}$.
It appears that only a part (which we will refer to as “transverse subspace”) of the eigenbasis is necessary to decompose $m^{\alpha}$ and $B^{\alpha}$. Furthermore, all eigenvalues of this subspace are $N_{r}$-times degenerate and only one vector from the linear space spanned by the $N_{r}$ corresponding eigenvectors enters the decomposition. The latter can be written as
$B^{\beta}_{\alpha a}=\sum_{\lambda}b^{\lambda}v^{\lambda\beta}_{\alpha a}=\sum_{\lambda}b^{\lambda}f^{\lambda}_{a}B^{\beta}_{\alpha a}.$ (21)
We observe that the respective eigenvectors of the transverse subspace have the same magnetization direction as the deforming transverse magnetic fields but different magnitudes locally in space given by the factors $f^{\lambda}_{a}$. Note that $\sum_{\lambda}b^{\lambda}f^{\lambda}_{a}=1$ at every point in space. The decomposition of the magnetization reads
$m^{\beta}_{\alpha a}=\sum_{\lambda}w^{\lambda}b^{\lambda}f^{\lambda}_{a}B^{\beta}_{\alpha a}.$ (22)
Eqs. (21) and (22) allow to write the transverse part of the ex-kernel as
$K^{\perp}_{a}=-{\mu_{\textrm{B}}}\frac{B^{\beta}_{\alpha a}}{m^{\beta}_{\alpha a}}=-{\mu_{\textrm{B}}}\frac{\sum_{\lambda}b^{\lambda}v^{\lambda\beta}_{\alpha a}}{\sum_{\lambda}w^{\lambda}b^{\lambda}v^{\lambda\beta}_{\alpha a}}=-{\mu_{\textrm{B}}}\frac{1}{\sum_{\lambda}w^{\lambda}b^{\lambda}f^{\lambda}_{a}}.$ (23)
One notes that the expression above is equivalent to Eq. (14) as $\sum_{\lambda}w^{\lambda}b^{\lambda}f^{\lambda}_{a}$ is the inverse ratio of the magnitudes of KS magnetic field and magnetization density which in the ALSDA point always in the same direction.
The discussion above allows to construct $K^{\perp}$ fulfilling the Goldstone sum rule exactly. Instead of computing $K^{\perp}$ from $m^{\mathrm{GS}}$ and $B^{\mathrm{GS}}$ using Eq. (14) one can find $b^{\lambda}$ from the the decomposition of one of the pairs $\left(m^{\alpha},B^{\alpha}\right)$} in the eigenvectors of the transverse subspace.
Here, we remark briefly that in our new scheme, with the improved spatial basis, providing the convergence with respect to the Brillouin zone and energy integrations is sufficient, this correction is hardly necessary as the Goldstone sum rule is fulfilled with accuracy of few meV for $q=0$ and the relative error of the pertinent degeneracy of eigevalues is about $10^{-5}$.
Let us conclude this section with the remark that the degeneracy $N_{r}=2$ is trivially given in the collinear case where the transverse part of the $q=0$ and $\omega=0$ susceptibility is block diagonal and isotropic
$\begin{pmatrix}\chi^{\mathrm{KS}}_{xaxb}&0\\ 0&\chi^{\mathrm{KS}}_{yayb}\end{pmatrix},\quad\chi^{\mathrm{KS}}_{yayb}=\chi^{\mathrm{KS}}_{xaxb}$ (24)
The corresponding transverse subspace is associated with the magnetization tilting around directions perpendicular to the ground state magnetization. (In our case, they lie in $x$-$y$ plane). In every spin channel independently, $v^{\lambda}_{a}$ are simply eigenvectors of the $\chi^{\mathrm{KS}}_{xaxb}$ block.
3 Implementation
3.1 Korringa-Kohn-Rostoker Green’s function method
We calculate the Korringa-Kohn-Rostoker (KKR) Green’s function (GF) for a NC ground state following the standard approach [Mankovsky2011, Simon2019]. First the single site scattering problem is solved yielding the corresponding regular and irregular solutions, $R_{\sigma L\sigma^{\prime}L^{\prime}}^{m}(r_{m},z)$ and $H_{\sigma L\sigma^{\prime}L^{\prime}}^{m}(r_{m},z)$ as well as the corresponding $t$-matrix. Following this, the backscattering operator taking into account the multiple electron-scattering in the lattice is found as
$\mathcal{G}_{\sigma L\sigma^{\prime}L^{\prime}}^{mn}(z)=g_{LL^{\prime}}^{mn}(z)\delta_{\sigma\sigma^{\prime}}+g_{LL_{1}}^{mk}(z)t_{\sigma L_{1}\sigma_{2}L_{2}}^{k}(z)\mathcal{G}_{\sigma_{2}L_{2}\sigma^{\prime}L^{\prime}}^{kn}(z)$ (25)
where $g_{LL^{\prime}}^{mn}(z)$ is the reference Green’s function. This allows to write down the KKR Green’s function as
$G_{\sigma\sigma^{\prime}}({\mathbf{x}},{\mathbf{x}}^{\prime},z)=$ $\Bigl[R_{\sigma L\sigma_{1}L1}^{m}(r_{m},z)\mathcal{G}_{\sigma_{1}L_{1}\sigma_{2}L_{2}}^{mn}(z)R_{\sigma_{2}L_{2}\sigma^{\prime}L^{\prime}}^{n}(r_{n},z){}$
$+$ $\sqrt{z}\delta_{mn}R_{\sigma L\sigma_{1}L_{1}}^{m}(r_{m}^{<},z)H_{\sigma_{1}L_{1}\sigma^{\prime}L^{\prime}}^{m}(r_{n}^{>},z)\Bigr]{\mathcal{Y}}_{L}(\hat{{\mathbf{r}}}_{m}){\mathcal{Y}}_{L^{\prime}}(\hat{{\mathbf{r}}}_{n})$ (26)
Einstein sum convention is used in equations 25 and 26. The KKR approach yields an all-electron picture of the Kohn-Sham system with the core states and atomic potentials found self-consistently. We assume the system to feature the discrete translational symmetry including the magnetization orientation as well and, thus, we do not deploy the generalized Bloch theorem for incommensurate spirals in the sense of Sandratskii [Sandratskii1991].
Our KKR scheme is capable of describing full scattering potentials, i.e., of arbitrary non-spherical shape, including varying the direction of magnetization within an atomic cell (intra-atomic NCity) [Papanikolaou2002]. To this end, the shape functions [Zeller1987] are used to define the Wigner-Seitz polyhedra [Stefanou1990] both when computing the magnetization and charge densities as well as when evaluating the overlap integrals in the susceptibility Dyson equation (2.2). However, the closed lattice structure of the systems under consideration allows to resort to the atomic sphere or full-charge density approximation. Furthermore, we observe that the intra-atomic NCity in the systems under consideration in this work is practically negligible [Sharma2007] and, thus, assume that the magnetization points along a certain common direction within the given atomic polyhedron $s$ defined by the angles $\vartheta_{s}$ and $\varphi_{s}$.
3.2 Evaluation of susceptibility
We turn now to the details of the evaluation of the density response functions. The complex plane integration scheme allowing to evaluate Eq. (7) is exposed in detail in [Buczek2011a]. It does not differ from the collinear case. The spatial $\left({\mathbf{x}},{\mathbf{x}^{\prime}}\right)$} dependence of $\chi$ is treated as follows. The discrete periodicity of the atomic system, if any, is used to diagonalize the susceptibility in the Bloch basis and cast it into the form of $\chi{\left({\mathbf{r}},{\mathbf{r}^{\prime}},{\mathbf{q}}\right)}$, where $\mathbf{r}$} and $\mathbf{r}^{\prime}$} belong to the primitive cell of the crystal and $\mathbf{q}$} resides in the first Brillouin zone $\Omega_{\textrm{BZ}}$ of the crystal. This Fourier transformation of the susceptibility is obtained from the convolution of two KKR Green’s function over $\Omega_{\textrm{BZ}}$. The spatial dependence of the susceptibility in the primitive cell is developed in the atomic centered coordinates (density indices $ij$ are suppressed for clarity) as
$\chi{\left({\mathbf{r}},{\mathbf{r}^{\prime}}\right)}=\sum_{LL^{\prime}}\chi^{ss^{\prime}}_{LL^{\prime}}{\left(r_{s},r_{s^{\prime}}\right)}{\mathcal{Y}}_{L}{\left(\hat{{\mathbf{r}}}_{s}\right)}{\mathcal{Y}}_{L^{\prime}}{\left(\hat{{\mathbf{r}}}_{s^{\prime}}\right)}$ (27)
Here, ${\mathbf{r}}_{s}$ should be understood as a reduced vector $\mathbf{r}$}, ${\mathbf{r}}_{s}\equiv{\mathbf{r}}-{\mathbf{s}}_{s}$, where ${\mathbf{s}}_{s}$ is the center of the atomic polyhedron containing $\mathbf{r}$}. The angular dependence of the spatial arguments in each cell is developed using spherical harmonics ${\mathcal{Y}}_{L}$ with index $L\equiv lm$ and argument $\hat{{\mathbf{r}}}_{s}\equiv{\mathbf{r}}_{s}/{\left|{\mathbf{r}}_{s}\right|}$. Note that in the case of full potential calculations, the product of two shape functions for the polyhedra $s$ and $s^{\prime}$ is incorporated into the definition of $\chi^{ss^{\prime}}_{LL^{\prime}}{\left(r_{s},r_{s^{\prime}}\right)}$. The radial dependence of $\chi{\left(r_{s},r_{s^{\prime}}\right)}$ can be developed in different ways. In our previous works, cf., e.g., [Buczek2011a], we took advantage of Chebyshev polynomials in order to achieve a rapid convergence with respect to the number of basis functions, as, typically, only around 12 were sufficient to fairly describe the dependence. In the current work, the function $\chi{\left(r_{s},r_{s^{\prime}}\right)}$ is given directly in the real space on a Gaussian mesh between 0 and the radius of the bounding atomic sphere $s$. This approach leads to somewhat increased memory consumption but greatly enhances the fulfilment of the Goldstone sum rule. Upon this step, the susceptibilities are represented as ordinary two-dimensional matrices $\chi_{\lambda\lambda^{\prime}}$ where the super-index $\lambda={\left(i,s,L,\rho\right)}$ comprises the density channel designation $i$, site $s$, angular quantum numbers $L\equiv lm$, and the radial index $\rho$. In this representation, Eq. (2.2) can be solved by the following direct matrix inversion [Staunton1999, Staunton2000, Thakor2003].
$\chi={\left({\sf I}-\chi_{\mathrm{KS}}{\left(K_{\mathrm{xc}}+v_{\textrm{C}}\right)}\right)}^{-1}\chi_{\mathrm{KS}}.$ (28)
One particular aspect of density response evaluation in NC systems turns out to become particularly algorithmically challenging compared to the collinear case, namely the evaluation of the trace over spin indices in Eq. (8). We recall that the non-local part of the KKR Green’s function consists of a product of three spin-matrices such that there emerges a contribution to $S_{ij}$ as follows
$c_{ij}=\operatorname{tr}\sigma^{i}\nu_{0}\nu_{1}\nu_{2}\sigma^{j}\nu_{3}\nu_{4}\nu_{5}.$ (29)
where the blocks $\nu_{m}^{\zeta}$ of each spin matrix
$\nu_{m}=\begin{pmatrix}\nu_{m}^{0}&\nu_{m}^{2}\\ \nu_{m}^{3}&\nu_{m}^{1}\end{pmatrix}$ (30)
are themselves matrices with the site, angular, and radial indices and $\operatorname{tr}$ stands for spin index trace. To be specific, when the non-local contribution to the susceptibility is concerned, $\nu_{1}$ and $\nu_{4}$ represent the backscattering operators and the remaining $\nu_{m}$ denote regular single site scattering solutions. While there is conceptually no difficulty in performing the matrix multiplication and the trace operation directly, this brute force approach to the calculation of the spin matrix elements features two major disadvantages which ultimately would lead to serious deterioration of the numerical performance. First, the blocks $\nu_{m}^{\zeta}$ can become sparse, e.g., concerning the angular indices when no full potential or spin-orbit corrections are taken into account. The direct matrix multiplication would then produce terms known to be zero from the onset. Second, a particular product of these blocks (out of 256 appearing in the trace)
$\prod_{m=0}^{5}\nu_{m}^{\zeta_{m}}$ (31)
can contribute to several out of 16 different pairs of density indices $ij$. If Eq. (29) were to be repeatedly evaluated for varying $i$ and $j$, this product defined by specific $\left\{\zeta_{0},\ldots,\zeta_{5}\right\}$} would be unnecessarily repeatedly evaluated multiple times as well. In order to manage this complexity, we developed a symbolic algebra scheme for computing the contribution of different products to the given $ij$ pair. Since the product in Eq. (31) involves matrices, it is necessary to keep track of the order of $\nu_{m}^{\zeta_{m}}$ in it. To this end, we devised a new method for handling noncommutative algebraic operations and implemented it using symbolic algebra language primitives of Mathematica [WolframResearch2025]. Subsequently, the symbolic results are used to automatically generate the FORTRAN code deployed in the high-performance computing of the traces. A more detailed exposition of the algorithm is given in Appendix B. This strategy feature several decisive advantages over the traditional manual implementation of scientific software. It combines the ease of symbolic manipulation with the numerical performance of low-level programming languages without the necessity of the tedious manual implementation in the latter. Furthermore, a thorough automatic validation of symbolic models is decisively easier than the one of the low level software.
4 Magnons of IrMn3
Let us apply our formalism to the TAF IrMn\$_3\$, in which the Mn atoms, exhibiting a finite magnetic moment, form a NC structure. More precisely, we take a look at the characteristic 111 plane of the system. Figure 1 shows the magnon dispersion and the inverse lifetimes (shown as bars designating the full with at half-maximum, FWHM, of the magnon peak) obtained from our LRTDDFT calculations. Three linearly dispersing magnon modes emerge in the long wave-length limit from three distinct Goldstone modes, as expected from the symmetry analysis outlined above. Their maximum energies at the edge of the Brillouin zone amounts to about 370meV. The modes are clearly Landau damped but remain well-defined for all momenta. In general, the damping increases with energy but differs strongly depending on the magnon polarization.
Figure 1: Magnon-spectrum within the 2D Brillouin zone, i.e. magnons propagating in the 111-lattice plane. Error bars inform about the FWHM of the magnon peak, i.e. the inverse lifetime. On the right, the 2D Fermi surface is shown together with the path through the Brillouin zone.
5 Summary and discussion
We presented a novel implementation of the ab initio linear response time-dependent density functional theory addressing spin excitations in non-collinear magnets based on the Korringa-Kohn-Rostoker Green’s function method. We elaborated the details of the numerical implementation, focusing on the managing the algorithmic complexity by means of symbolic computational algebra, exposed the generalization of the adiabatic local spin density approximation to the exchange-correlation kernel to the non-collinear case. The emergence of multiple Nambu-Goldstone bosons in non-collinear susceptibility calculations is elaborated on formally and from the numerical convergence point of view, linked to the zero-energy magnetic deformation eigenvectors. With the aid of the formalism, we analysed the dispersion and Landau damping of magnons in kagome TAF Mn3Ir. We hope that this work will spark further developments in the non-collinear magnonics.
Acknowledgments. We thank Leonid M. Sandratskii, László Szunyogh, Patrick Perndorfer, and Sebastian Paischer for fruitful discussions and Igor Maznichenko for his advice and helpful remarks. P.B. and A.E. acknowledge the funding, respectively, by Österreichischer Fonds zur Förderung der Wissenschaftlichen Forschung (FWF) under grant I 5384 and by Deutsche Forschungsgemeinschaft (DFG) under grant BU 4062/1-1.
Appendix A Källén-Lehmann spectral representation and the
fluctuation-dissipation theorem
The numerical evaluation of the density response functions is a notoriously difficult undertaking and the following relations [Nyquist1928, Callen1951, Fetter1971, Kubo1966, Abrikosov1975, Giuliani2005] appeared useful in the validation of our computer implementation and the interpretation of results. The relations given below are completely general and apply even to the cases of NCity and non-centrosymmetric crystals with relativistic effects taken into account.
In order to simply the notation, we consider the response function between two operators labelled by $\alpha$. To be specific, in our case
$\alpha={\left(i{\mathbf{x}}\right)}$ (32)
but the following consideration apply to any Hermitian operators being time-independent in the Schrödinger picture. In order to study the response functions in periodic solids, we will nevertheless assume that the operators are ascribed to a particular primitive cell $\mathbf{R}$}. The retarded response function $\chi^{\mathcal{R}}$ is now defined as
$\chi_{\alpha\beta}^{\mathcal{R}}{\left({\mathbf{R}}-{\mathbf{R}}^{\prime},t-t^{\prime}\right)}\equiv-{{\mathfrak{i}}}{\theta}{\left(t-t^{\prime}\right)}{\left\langle{{\left[\hat{o}_{\alpha{\mathbf{R}}}{\left(t\right)},\hat{o}_{\beta{\mathbf{R}}^{\prime}}{\left(t^{\prime}\right)}\right]}}\right\rangle}.$ (33)
Since the function connects the change of the observable physical (real) observable represented by $\hat{o}_{\alpha{\mathbf{R}}}$ due to the disturbing physical field $\hat{o}_{\beta{\mathbf{R}}^{\prime}}$ and as such it is a real scalar, $\chi_{\alpha\beta}^{\mathcal{R}}{\left({\mathbf{R}},t\right)}\in{\mathbb{R}}$. Consequently, the time and lattice Fourier transformation (cf. Eq. (2)) obeys the conjugation symmetry
$\chi_{\alpha\beta}^{\mathcal{R}}{\left(-{\mathbf{q}},-\omega\right)}^{\ast}=\chi_{\alpha\beta}^{\mathcal{R}}{\left({\mathbf{q}},\omega\right)}.$ (34)
We remark that the above formula applies also to the case when the response function is evaluated not directly above the real axis ($\gamma>0$ in Eq. (7))
The Fourier-transformed susceptibility is conveniently expressed in the Källén-Lehmann spectral representation as
$\chi_{\alpha\beta}^{\mathcal{R}}{\left(\omega,{\mathbf{q}}\right)}=\sum_{{\mathbf{R}}\in\mathcal{L}}{\mathfrak{e}^{-{{\mathfrak{i}}}{\mathbf{q}}\cdot{\mathbf{R}}}}\sum_{sf}\frac{{\mathfrak{e}^{-E_{s}/({k_{\textrm{B}}}T)}}}{Z}{\left(\frac{\braket{s|\hat{o}_{\alpha{\mathbf{R}}}|f}\braket{f|\hat{o}_{\beta{\mathbf{0}}}|s}}{\omega-{\left(E_{f}-E_{s}\right)}+{{\mathfrak{i}}}{0^{+}}}-\frac{\braket{s|\hat{o}_{\beta{\mathbf{0}}}|f}\braket{f|\hat{o}_{\alpha{\mathbf{R}}}|s}}{\omega+{\left(E_{f}-E_{s}\right)}+{{\mathfrak{i}}}{0^{+}}}\right)},$ (35)
[…]
One quantifies the natural (spontaneous) fluctuations of the ground state observables using the following correlation function
${\mathfrak{C}}_{\alpha\beta}{\left({\mathbf{R}},t\right)}\equiv{\left\langle\hat{o}_{\alpha{\mathbf{R}}}{\left(t\right)}\hat{o}_{\beta{\mathbf{0}}}\right\rangle}$ (36)
Fourier-transformed as
${\mathfrak{C}}_{\alpha\beta}{\left({\mathbf{q}},\omega\right)}=2\pi\sum_{{\mathbf{R}}\in\mathcal{L}}{\mathfrak{e}^{-{{\mathfrak{i}}}{\mathbf{q}}\cdot{\mathbf{R}}}}\sum_{sf}\frac{{\mathfrak{e}^{-E_{s}/({k_{\textrm{B}}}T)}}}{Z}\braket{s|\hat{o}_{\alpha{\mathbf{R}}}|f}\braket{f|\hat{o}_{\beta{\mathbf{0}}}|s}\delta{\left(\omega-{\left(E_{f}-E_{s}\right)}\right)}.$ (37)
Using the loss matrix
$L_{\alpha\beta}{\left({\mathbf{q}},\omega\right)}\equiv\frac{1}{2{{\mathfrak{i}}}}{\left(\chi^{\mathcal{R}}_{\alpha\beta}{\left({\mathbf{q}},\omega\right)}-\chi^{\mathcal{R}}_{\beta\alpha}{\left({\mathbf{q}},\omega\right)}^{\ast}\right)}$ (38)
the correlation function is written as
${\mathfrak{C}}_{\alpha\beta}{\left({\mathbf{q}},\omega\right)}=-\frac{2}{1-{\mathfrak{e}^{-\omega/({k_{\textrm{B}}}T)}}}L_{\alpha\beta}{\left({\mathbf{q}},\omega\right)}.$ (39)
$\mathfrak{C}$}, similar to $L$, is a Hermitian matrix
${\mathfrak{C}}_{\alpha\beta}{\left({\mathbf{q}},\omega\right)}={\mathfrak{C}}_{\beta\alpha}{\left({\mathbf{q}},\omega\right)}^{\ast}.$ (40)
One also proves that
$L_{\alpha\beta}{\left({\mathbf{q}},\omega=0\right)}=0$ (41)
which in turn implies that $\chi_{\alpha\beta}^{\mathcal{R}}{\left({\mathbf{q}},\omega=0\right)}$ is a Hermitian and $\chi_{\alpha\beta}^{\mathcal{R}}{\left({\mathbf{q}}={\mathbf{0}},\omega=0\right)}$ a symmetric matrix.
Let us assume that our system is weakly and mono-chromatically perturbed by a Hamiltonian of the following form
$\hat{H}{\left(t\right)}=\sum_{\alpha}f_{\alpha}\hat{o}_{\alpha{\mathbf{R}}}\cos{\left({\mathbf{q}}\cdot{\mathbf{R}}-\omega t-\phi\right)}.$ (42)
It is given in the Schrödinger picture. $f_{\alpha}$ are real scalars representing the potentials coupling to operators $\hat{o}_{\alpha}$. The power absorbed from this perturbing field reads
$P=-\frac{\omega}{2}\sum_{{\mathbf{R}},{\mathbf{R}}\in\mathcal{L}}{\mathfrak{e}^{-{{\mathfrak{i}}}{\mathbf{q}}\cdot{\left({\mathbf{R}}-{\mathbf{R}}^{\prime}\right)}}}\sum_{\alpha\beta}f_{\alpha}L_{\alpha{\mathbf{R}}\beta{\mathbf{R}}^{\prime}}{\left(\omega\right)}f_{\beta}=-N\frac{\omega}{2}\sum_{\alpha\beta}f_{\alpha}L_{\alpha\beta}{\left({\mathbf{q}},\omega\right)}f_{\beta}.$ (43)
In the above equation, $N$ stands for the number of primitive cells of the crystal under perturbation. As expected, $P$ is an extensive quantity. We observe that the absorption is given by the eigenvalues and eigenvectors of $L$. Since the absorption occurs due to the generation of the system’s excited states, the eigensystem of $L$ yields them directly.
The fluctuation-dissipation theorem reflects the observation that both the spontaneous fluctuations of the ground state observables (Eq. (39)) as well as the absorption of power from external driving field (Eq. (43)) are given by the anti-Hermitian part of the response function $L$.
Furthermore, ${\mathfrak{C}}_{\alpha\beta}{\left({\mathbf{q}},\omega\right)}$ allows to write down the Kramers-Kronig relation for the response function [Kronig1926] as
$\chi_{\alpha\beta}^{\mathcal{R}}{\left({\mathbf{q}},\omega\right)}=$ $\int\frac{d\omega^{\prime}}{2\pi}\frac{{\mathfrak{C}}_{\alpha\beta}{\left({\mathbf{q}},\omega^{\prime}\right)}}{\omega-\omega^{\prime}+{{\mathfrak{i}}}{0^{+}}}-\int\frac{d\omega^{\prime}}{2\pi}\frac{{\mathfrak{C}}_{\beta\alpha}{\left(-{\mathbf{q}},\omega^{\prime}\right)}}{\omega+\omega^{\prime}+{{\mathfrak{i}}}{0^{+}}}.$ (44)
The above allows to show that
${\mathfrak{C}}_{\alpha\beta}{\left({\mathbf{q}},\omega\right)}-{\mathfrak{C}}_{\beta\alpha}{\left(-{\mathbf{q}},-\omega\right)}=-2L_{\alpha\beta}{\left({\mathbf{q}},\omega\right)}.$ (45)
Finally, for completeness, we write down equivalent expressions for the advanced ($\chi^{\mathcal{A}}$), causal ($\chi$), and thermal ($\chi^{\mathcal{T}}$) response functions
$\chi_{\alpha\beta}^{\mathcal{A}}{\left({\mathbf{R}}-{\mathbf{R}}^{\prime},t-t^{\prime}\right)}$ $\equiv{{\mathfrak{i}}}{\theta}{\left(t^{\prime}-t\right)}{\left\langle{{\left[\hat{o}_{\alpha{\mathbf{R}}}{\left(t\right)},\hat{o}_{\beta{\mathbf{R}}^{\prime}}{\left(t^{\prime}\right)}\right]}}\right\rangle}.$ (46)
$\chi_{\alpha\beta}{\left({\mathbf{R}}-{\mathbf{R}}^{\prime},t-t^{\prime}\right)}$ $\equiv-{{\mathfrak{i}}}{\left\langle{\operatorname{T}{\left[\hat{o}_{\alpha{\mathbf{R}}}{\left(t\right)}\hat{o}_{\beta{\mathbf{R}}^{\prime}}{\left(t^{\prime}\right)}\right]}}\right\rangle}$ (47)
$\chi_{\alpha\beta}^{\mathcal{T}}{\left({\mathbf{R}}-{\mathbf{R}}^{\prime},\tau-\tau^{\prime}\right)}$ $\equiv-{\left\langle{\operatorname{T}{\left[\hat{o}_{\alpha{\mathbf{R}}}{\left(\tau\right)}\hat{o}_{\beta{\mathbf{R}}^{\prime}}{\left(\tau^{\prime}\right)}\right]}}\right\rangle}.$ (48)
The corresponding spectral representations for the first two read
$\chi_{\alpha\beta}^{\mathcal{A}}{\left({\mathbf{q}},\omega\right)}=$ $\int\frac{d\omega^{\prime}}{2\pi}\frac{{\mathfrak{C}}_{\alpha\beta}{\left({\mathbf{q}},\omega^{\prime}\right)}}{\omega-\omega^{\prime}-{{\mathfrak{i}}}{0^{+}}}-\int\frac{d\omega^{\prime}}{2\pi}\frac{{\mathfrak{C}}_{\beta\alpha}{\left(-{\mathbf{q}},\omega^{\prime}\right)}}{\omega+\omega^{\prime}-{{\mathfrak{i}}}{0^{+}}},$ (49)
$\chi_{\alpha\beta}{\left({\mathbf{q}},\omega\right)}=$ $\int\frac{d\omega^{\prime}}{2\pi}\frac{{\mathfrak{C}}_{\alpha\beta}{\left({\mathbf{q}},\omega^{\prime}\right)}}{\omega-\omega^{\prime}+{{\mathfrak{i}}}{0^{+}}}-\int\frac{d\omega^{\prime}}{2\pi}\frac{{\mathfrak{C}}_{\beta\alpha}{\left(-{\mathbf{q}},\omega^{\prime}\right)}}{\omega+\omega^{\prime}-{{\mathfrak{i}}}{0^{+}}}.$ (50)
On the other hand, the Matsubara representation of $\chi^{\mathcal{T}}$ contains only bosonic frequencies ${\omega_{m}^{b}}=2m\pi{k_{\textrm{B}}}T$, and its Fourier transformation and the corresponding spectral representation reads
$\chi_{\alpha\beta}^{\mathcal{T}}{\left({\mathbf{q}},{\omega_{m}^{b}}\right)}$ $\equiv\int_{0}^{1/{\left({k_{\textrm{B}}}T\right)}}\chi_{\alpha\beta}^{\mathcal{T}}{\left({\mathbf{q}},\tau\right)}{\mathfrak{e}^{{{\mathfrak{i}}}{\omega_{m}^{b}}\tau}}$ (51)
$=\int\frac{d\omega}{2\pi}\frac{{\mathfrak{C}}_{\alpha\beta}{\left({\mathbf{q}},\omega\right)}}{{{\mathfrak{i}}}{\omega_{m}^{b}}-\omega}-\int\frac{d\omega}{2\pi}\frac{{\mathfrak{C}}_{\beta\alpha}{\left(-{\mathbf{q}},\omega\right)}}{{{\mathfrak{i}}}{\omega_{m}^{b}}+\omega}.$ (52)
Appendix B The symbolic handling of the spin matrix traces
The evaluation of the dynamical susceptibility based on the KKR Green’s functions involves performing the spin index trace of the product of multiple matrices in the typical form of the one in Eq. (29). To this end, we devised a method for handling noncommutative algebraic operations motivated by the Gödel numbering developed by him in the context of the proof of his incompleteness theorem [Goedel1931]. The specific implementation is formulated in the symbolic algebra language primitives in Mathematica [WolframResearch2025]. The matrices $\nu_{m}$ are represented symbolically as
$\nu_{m}=\begin{pmatrix}x^{0\cdot 4^{m}}&x^{2\cdot 4^{m}}\\ x^{3\cdot 4^{m}}&x^{1\cdot 4^{m}}\end{pmatrix}$, (53)
where $\cdot$ denotes ordinary multiplication. The matrix multiplication of such matrices
$\prod_{m=0,1,\ldots}\nu_{m}$ (54)
and subsequent trace taking generates polynomials of a form allowing for an automatic identification of the original position of a specific factor in the above product. It relies on the fact that the power of $x$ expressed in quaternary numeral system can be directly interpreted as the sought-after order.
As a simpler example, we analyse the charge and spin density calculations from the KKR Green’s function having the form of
$\rho^{i}=\operatorname{tr}\sigma^{i}\nu_{0}\nu_{1}\nu_{2}.$ (55)
When the form given by Eq. (53) is inserted, one obtains the following column vector for the density channels
$\begin{pmatrix}\rho^{0}\\ \rho^{x}\\ \rho^{y}\\ \rho^{z}\end{pmatrix}=\begin{pmatrix}x^{56}+x^{54}+x^{45}+x^{35}+x^{27}+x^{21}+x^{14}+1\\ x^{59}+x^{53}+x^{46}+x^{32}+x^{24}+x^{22}+x^{13}+x^{3}\\ -{{\mathfrak{i}}}{\left(x^{59}+x^{53}-x^{46}-x^{32}-x^{24}-x^{22}+x^{13}+x^{3}\right)}\\ x^{56}+x^{54}-x^{45}-x^{35}-x^{27}-x^{21}+x^{14}+1\end{pmatrix}$ (56)
To be specific, let us consider the $x^{24}$ term. In the quaternary representation $24_{10}=120_{4}$ and corresponds element $1$ of $\nu_{2}$ ($\nu_{2}^{22}$), element $2$ of $\nu_{1}$ ($\nu_{1}^{12}$), and element 0 of $\nu_{0}$ ($\nu_{0}^{11}$). We see that this product appears in the expression for $\rho^{x}$ and $\rho^{y}$ spin density with the coefficients of $1$ and $\mathfrak{i}$}}, respectively. In other words, in this specific calculation, the product
$\nu_{0}^{11}\nu_{1}^{12}\nu_{2}^{22}$ (57)
is computed only once but used twice, in the sum yielding $\rho^{x}$ (contribution $\nu_{0}^{11}\nu_{1}^{12}\nu_{2}^{22}$) and $\rho^{y}$ (contribution ${{\mathfrak{i}}}\nu_{0}^{11}\nu_{1}^{12}\nu_{2}^{22}$). The generalization to the case of products containing $6$ spin matrices, cf. Eq. (29), follows the same pattern. In this case, one obtains $256$ different $6$-products of $\nu$’s, each used $4$ times in different $ij$-pairs of spin channels. With the determined contributions and coefficients at hand, the automatic generation of the FORTRAN computer code is facilitated.